# Unraveling the Enigma of Double Descent: An In-depth Analysis through the Lens of Learned Feature Space

### v1

@@ Abstract
Double descent presents a counter-intuitive aspect within the machine learning domain, and researchers have observed its manifestation in various models and tasks. While some theoretical explanations have been proposed for this phenomenon in specific contexts, an accepted theory to account for its occurrence in deep learning remains yet to be established. In this study, we revisit the phenomenon of double descent and demonstrate that its occurrence is strongly influenced by the presence of noisy data. Through conducting a comprehensive analysis of the feature space of learned representations, we unveil that double descent arises in imperfect models trained with noisy data. We argue that double descent is a consequence of the model first learning the noisy data until interpolation and then adding implicit regularization via over-parameterization acquiring therefore capability to separate the information from the noise. We postulate that double descent should never occur in well-regularized models.
@@ 1 Introduction
As a well-known property of machine learning, bias-variance trade-off suggested that the variance of the parameter estimated across samples can be reduced by increasing the bias in the estimated parameters Geman et al. (1992). When a model is under-parameterised, there exists a combination of high bias and low variance. Under this circumstance, the model becomes trapped in under-fitting, signifying its inability to grasp the fundamental underlying structures present in the data. When a model is over-parameterised, there exists a combination of low bias and high variance. This is usually due to the model over-fitting noisy and unrepresentative data in the training set. Traditional statistical intuition suggests a difficulty in reducing bias and variance simultaneously and advises finding a “sweet-spot” between underfitting and overfitting. This conjugate property prevents machine learning estimators from generalizing well beyond their dataset Hastie et al. (2009). However, modern machine learning practices have achieved remarkable success with the training of highly-parameterized sophisticated models that can precisely fit the training data (‘interpolate’ the training set) while attaining impressive performance on unseen test data.
@@
The double descent phenomenon is prevalent across various machine learning models. In the context of linear models, a series of empirical and theoretical studies have emerged to analyze this phenomenon Loog et al. (2020). In deep learning, a number of researchers have offered additional evidence supporting the presence and prevalence of the double descent phenomenon across a broader range of deep learning models, including CNNs, ResNets, and Transformers, in diverse experimental setups. Nakkiran et al. (2021) also reveals that this phenomenon is most pronounced when dealing with label noise in the training set, while can manifest even without label noise. Despite the widely accepted empirical observation of the double descent phenomenon within the domain of deep learning, comprehending the underlying mechanism of this occurrence in deep neural networks continues to pose a significant question.
In this paper, we highlight the crucial role that noise plays in shaping the behavior of deep learning models through a comprehensive exploration and interpretation of the interplay among model width, noisy training data and the learned feature space. We refer the learned feature space as the space constructed from learned representations within fully-trained neural networks, which is mapped from the training data and directly utilized as supporting evidence in classification tasks. We unveil that as model width expands, the model becomes capable of segregating inaccurately labeled data within the acquired feature space. This finding motivates us to propose that the ability of the model to adapt and recognize noisy data has a substantial impact on its overall generalization performance and contributes to the emergence of the double descent phenomenon. Indeed, our study suggests that this phenomenon is a consequence of models learning how to separate the noise from the information in the training set, and we offer valuable perspectives into the performance and constraints of deep learning models, thereby opening avenues for enhanced approaches to address noisy data and amplify model generalization capabilities.
The main contributions of this study are summarized as follows. First, we extensively examine model-wise deep double descent for a wide range of classification tasks and models. We proposed that there exists a robust correlation between the distribution of noisy labeled training data in the learned feature space and the double descent phenomenon. Through the $k$-nearest neighbour algorithm and T-Distributed Stochastic Neighbor Embedding (t-SNE) visualization, we provide evidence that within highly extensive models, noisy labeled data is effectively isolated among their true classmates, shedding light on the mechanism by which over-parameterization enhances generalization.
@@ 2 Related Work
The initial proposal of the double descent phenomenon as a general concept was made by Belkin et al. (2019) for increasing model size (model-wise). However, Nakkiran et al. (2021) have also shown a similar trend w.r.t. size of training dataset (sample-wise) and training time (epoch-wise). While the double descent phenomenon is widely discussed in these alternative perspectives, we focus our attention towards the original model-wise phenomenon in this work, briefed as ‘double descent’ in the subsequent discussion.
@@
Double Descent has been shown as a robust phenomenon that occurs over diverse tasks, architectures, and optimization techniques. Subsequently, a substantial volume of research has been conducted to investigate the double descent phenomenon regarding its mechanism. A stream of theoretical literature have emerged, delving into the establishment of generalization bounds and conducting asymptotic analyses, particularly focused on linear functions (Advani et al., 2020; Belkin et al., 2020; Bartlett et al., 2020; Muthukumar et al., 2021; Mei & Montanari, 2022; Hastie et al., 2022). Hastie et al. (2022) highlight that when the linear regression model is misspecified, the best generalization performance can occur in the over-parameterized regime. Mei & Montanari (2022) further concluded that the occurrence of the double descent phenomenon can be attributed to model misspecification, which arises when there is a mismatch between the model structure and the model family. While people believe the same intuition extends to deep learning as well, nevertheless, the underlying mechanism still eludes a comprehensive understanding of ‘deep’ double descent.
@@
The prevailing literature that explores the double descent phenomenon in deep neural networks learning provide a range of explanations: bias-variance decomposition Yang et al. (2020); samples to parameters ratio Belkin et al. (2020); Nakkiran et al. (2021); decision boundaries Somepalli et al. (2022); and sharpless of interpolation Gamba et al. (2022). While offering an extensive array of experimental findings that replicate the double descent phenomenon, Nakkiran et al. (2021) reached the conclusion that the emergence of double descent exhibits a robust correlation with the inclusion of noisy labels in the training dataset. Somepalli et al. (2022) further studied the decision boundaries of random data points and established a relationship between the fragmentation of class regions and the double descent phenomenon. Additionally, Teague (2022) derives an interesting intuition for implicit regularization from over-parameterization relating it to the geometrical properties of volumes of geometric figures in hyperspaces. Chaudhary et al. (2023) revisited the phenomenon and proposed that double descent may not inherently characterize deep learning.
@@
In contrast to existing studies on double descent, our approach distinguishes itself by exploring the phenomenon through the lens of the learned feature space generated by trained models. The most related works to ours are the concurrent ones by Gamba et al. (2022), which measure sharpness at clean and noisily-labeled data points and empirically showed that smooth interpolation emerging both for large over-parameterized networks and trained large models confidently predicts the noisy training targets over large volumes around each training point. While their study also entailed an analysis involving separate evaluations of clean and noisily-labeled data points, we employed the $k$-nearest neighbor algorithm to infer the interrelationship between these two categories of training data in the learned feature space and demonstrated that for over-parameterized models, noisy labeled data tends to be classified alongside its genuine counterparts. Our approach suggests that the double descent phenomenon is a consequence of deep learning models learning how to separate the noise from the information in the training set.
@@ 3 Methodology
We first assess the contribution of noisy labeled data to the double descent phenomenon. Our hypothesis is that the emergence of the double descent phenomenon can be attributed to the progressively over-parameterized models effectively isolating noisy data within the training set, thus diminishing the influence of interpolating these noisy data points. Since this proposition is inspired by Nakkiran et al. (2021)’s empirical conclusion, we initiate our experimentation by replicating the double descent phenomenon as per the experimental framework outlined by this paper. This approach offers a comprehensively manageable setup and establishes a comparative analysis of the phenomenon across diverse tasks and optimization techniques. With increasing ratio of label noise introduced to the training dataset, the test error peak at the interpolation threshold becomes evident, signifying a definite correlation with the pronounced double descent phenomenon. We then used the $k$-nearest neighbour algorithm to infer the relative locations of clean and noisy labelled data in the learned feature space. This methodology allows us to delineate the prediction strategy through an indirect approach and subsequently compare our findings with the overall generalization performance of these pre-trained models.
@@ 3.1 Experiment Setup
  • Standard Convolutional Networks (CNNs): We consider a family of standard convolutional neural networks formed by 4 convolutional stages of controlled base width [$k$, $2k$, $4k$, $8k$], for $k$ in the range of [1, 64], along with a fully connected layer as the classifier. The MaxPool is [2, 2, 2, 4]. For all the convolution layers, the kernel size = 3, stride = 1 and padding = 1. This architecture implementation is adopted from Nakkiran et al. (2021).
@@
All networks are trained using the Stochastic Gradient Descent (SGD) optimizer with Cross-Entropy Loss, zero momentum, and decreasing learning rate. A batch size of 128 is employed, and no explicit data augmentation or dropout techniques are utilized in our experimental setup. Utilizing a sequence of models trained using a uniform training scheme showcases the double descent phenomenon across different models. We refer our reader to Appendix A for further information regarding our experiment setup and scaling of model parameterization with layer width k.
@@ 3.2 Introduction of Label Noise
Our analysis revolves around the introduction of noisy labels. In our experimental setup, label noise with probability $p$ signifies that during training, samples are assigned a uniformly random label with probability $p$, while possessing the correct label with probability $(1-p)$ otherwise. We interpret a neural network with input variable $x\in R^{d}$ and learn-able parameter $\theta$, incorporating all weights and biases. The train dataset is defined as a set of training points $(x_{n},Y_{n})$, for $n=[1,...,N]$. In a training dataset containing $N$ data points, there exist $(1-p)N$ clean data points and $pN$ noisily labeled data points.
@@
It is important to note that the label noise is sampled only once and not per epoch, and this original and random label information is stored prior to each experiment. Each experiment is replicated multiple times, and the final result is obtained by averaging the outcomes.
3.3 Learned Feature Space Detection Strategy
@@
Our primary research inquiry aims to gain insights into the diverse strategies employed by various deep neural networks in interpolating these noisy labeled data across varying model capacities in classification tasks. Thus, our investigation centered on the feature space of learned representations, typically denoting the feed-forward features $f_{n}$, for $n=[1,...,N]$ of the penultimate layer prior to classification. Acknowledging the difficultries of inferring high-demensional feature spaces, our approach mainly entails inferring the relative positioning of training data points. This involves establishing a connection between the hidden features of noisy data points in the learned feature space and their $k$-nearest neighbors within the clean training dataset.
In order to compare the true label of a data point, which is assigned a random label (noisy labeled data), with the label of its $k$-nearest neighbors in the learned feature space; we calculate the prediction accuracy ${\rm Kp}$ of noisy labelled data of the original clean labels for each noisy labeled data point matching the label prediction of its $k$-nearest neighbors in the learned feature space:
${\rm Kp}=\frac{\sum_{i}Y_{i}\land\hat{Y}_{i}}{pN},$ (1)
where $\hat{Y}_{n}$ is the KNN prediction, and where the sum is over the data either the train set or the test set for the in-sample or out-of-sample computations. This prediction accuracy ${\rm Kp}$ is able to reflect the neural architecture’s interpolation strategy on these noisy labelled data. In addition to evaluating the value of ${\rm Kp}$, we also utilize T-Distributed Stochastic Neighbor Embedding (t-SNE) to provide a visual representation of the learned features in the feature space. While this visualization may not accurately represent the intricate inherent structure of high-dimensional features, it aims to gain insights into the clustering and distribution of data points, thereby enhancing our understanding of the model’s internal representations.
@@
In the next section, we present our empirical study and results of the learned feature space for trained neural networks in relationship to double descent.

### v2

@@ Abstract
Double descent presents a counter-intuitive aspect within the machine learning domain, and researchers have observed its manifestation in various models and tasks. While some theoretical explanations have been proposed for this phenomenon in specific contexts, an accepted theory for its occurrence in deep learning remains yet to be established. In this study, we revisit the phenomenon of double descent and demonstrate that its occurrence is strongly influenced by the presence of noisy data. Through conducting a comprehensive analysis of the feature space of learned representations, we unveil that double descent arises in imperfect models trained with noisy data. We argue that double descent is a consequence of the model first learning the noisy data until interpolation and then adding implicit regularization via over-parameterization acquiring therefore capability to separate the information from the noise.
@@ 1 Introduction
As a well-known property of machine learning, the bias-variance trade-off suggested that the variance of the parameter estimated across samples can be reduced by increasing the bias in the estimated parameters (Geman et al., 1992). When a model is under-parameterized, there exists a combination of high bias and low variance. Under this circumstance, the model becomes trapped in under-fitting, signifying its inability to grasp the fundamental underlying structures present in the data. When a model is over-parameterized, there exists a combination of low bias and high variance. This is usually due to the model over-fitting noisy and unrepresentative data in the training set. Traditional statistical intuition suggests a difficulty in reducing bias and variance simultaneously and advises finding a ‘sweet spot’ between underfitting and overfitting. This conjugate property prevents machine learning estimators from generalizing well beyond their dataset (Hastie et al., 2009). However, modern machine learning practices have achieved remarkable success with the training of highly parameterized sophisticated models that can precisely fit the training data (‘interpolate’ the training set) while attaining impressive performance on unseen test data.
@@
The double descent phenomenon is prevalent across various machine learning models. In the context of linear models, a series of empirical and theoretical studies have emerged to analyze this phenomenon (Loog et al., 2020). In deep learning, a number of researchers have offered additional evidence supporting the presence and prevalence of the double descent phenomenon across a broader range of deep learning models, including CNNs, ResNets, and Transformers, in diverse experimental setups. Nakkiran et al. (2021) also reveals that this phenomenon is most pronounced when dealing with label noise in the training set, while can manifest even without label noise. Despite the widely accepted empirical observation of the double descent phenomenon within the domain of deep learning, comprehending the underlying mechanism of this occurrence in deep neural networks continues to pose a significant question.
It is well-known that deep neural networks accomplish end-to-end learning through integrating feature learning with classifier training (Zhou, 2021). A neural network tailored for classification tasks can be divided into two components. The initial segment is dedicated to feature space transformation, specifically the conversion of the original feature space represented by the input layer to the final feature space represented by the final representation layer, in which a classifier is constructed. We refer to this final feature space as the learned feature space, which is shaped by the acquired representations within fully-trained neural networks. In this paper, we highlight the crucial role that noise plays in shaping the behavior of deep learning models through a comprehensive exploration and interpretation of the interplay among model width, noisy training data, and the learned feature space. We demonstrated that over-parameterized models when contrasted with under-parameterized ones, interpolate more correct training data around noisy ones of matching classes. This phenomenon demonstrates a correlation with the double descent phenomenon, and we propose that this is attributed to models acquiring the ability to effectively ‘isolate’ noise from the information in the training dataset. We posit that this phenomenon has not been previously documented and could serve to elucidate the internal mechanism of double descent.
The main contributions of this study are summarized as follows. We proposed that there exists a robust correlation between the neural network’s interpolation strategy of noisy data and the double descent phenomenon. Utilizing the $k$-nearest neighbor algorithm, we indicate that this phenomenon results from broader architectures genuinely interpolating noisy samples among correct samples of the same original class, showcasing an enhanced capability to isolate noise information. We believe our investigation is able to shed light on the mechanism by which over-parameterization enhances generalization.
@@ 2 Related Work
The initial proposal of the double descent phenomenon as a general concept was made by Belkin et al. (2019) for increasing model size (model-wise). However, Nakkiran et al. (2021) have also shown a similar trend w.r.t. size of the training dataset (sample-wise) and training time (epoch-wise). The research further reports observation of all forms of double descent most strongly in settings with label noise in the train set. While the double descent phenomenon is widely discussed in these alternative perspectives, we focus our attention on the original model-wise phenomenon in this work, briefed as ‘double descent’ in the subsequent discussion.
@@
Double Descent has been shown as a robust phenomenon that occurs over diverse tasks, architectures, and optimization techniques. Subsequently, a substantial volume of research has been conducted to investigate the double descent phenomenon regarding its mechanism. A stream of theoretical literature has emerged, delving into the establishment of generalization bounds and conducting asymptotic analyses, particularly focused on linear functions (Advani et al., 2020; Belkin et al., 2020; Bartlett et al., 2020; Muthukumar et al., 2021; Mei & Montanari, 2022; Hastie et al., 2022). Hastie et al. (2022) highlight that when the linear regression model is misspecified, the best generalization performance can occur in the over-parameterized regime. Mei & Montanari (2022) further concluded that the occurrence of the double descent phenomenon can be attributed to model misspecification, which arises when there is a mismatch between the model structure and the model family. While people believe the same intuition extends to deep learning as well, nevertheless, the underlying mechanism still eludes a comprehensive understanding of ‘deep’ double descent.
@@
The prevailing literature that explores the double descent phenomenon in deep neural networks learning provides a range of explanations: bias-variance decomposition (Yang et al., 2020); samples to parameters ratio (Belkin et al., 2020; Nakkiran et al., 2021); decision boundaries (Somepalli et al., 2022); and sharpless of interpolation (Gamba et al., 2022). While offering an extensive array of experimental findings that replicate the double descent phenomenon, Nakkiran et al. (2021) concluded that the emergence of double descent exhibits a robust correlation with the inclusion of noisy labels in the training dataset. Somepalli et al. (2022) further studied the decision boundaries of random data points and established a relationship between the fragmentation of class regions and the double descent phenomenon. Recently, Chaudhary et al. (2023) revisited the phenomenon and proposed that double descent may not inherently characterize deep learning.
Beyond the thorough examinations of double descent, an emerging theory of over-parameterized machine learning (TOPML) is growing, seeking to elucidate the second ‘descent’ in the double descent phenomenon (Dar et al., 2021). The inquiry posed by Zhou (2021) revolves around the puzzling aspect of why over-parameterized models do not succumb to overfitting and focus on the feature space transformation aspect. Chang et al. (2021) analytically identify regimes when training over-parameterized networks and then pruning is preferable to training small models directly. Teague (2022) intriguingly explores the concept of implicit regularization arising from over-parameterization and draws a connection between the generalization capacity and the geometric properties of volumes within hyperspaces. The work by Liu et al. (2022) introduces a robust training method for over-parameterized deep networks in classification tasks, specifically when training labels are corrupted. Although our research shares a thematic connection, our study concentrates on the implicit mechanisms through which over-parameterized models acquire the ability to discern and handle label noise. Consequently, our focus differs from their research.
@@
In contrast to existing studies, our approach distinguishes itself by exploring the phenomenon through the lens of the learned feature space generated by trained models. The most related works to ours are the concurrent ones by Gamba et al. (2022), which measure sharpness at clean and noisily-labeled data points and empirically showed that smooth interpolation emerging both for large over-parameterized networks and trained large models confidently predicts the noisy training targets over large volumes around each training point. While their study also entailed an analysis involving separate evaluations of clean and noisily labeled data points, we employed the $k$-nearest neighbor algorithm to infer the inter-relationship between these two categories of training data in the learned feature space and demonstrated that for over-parameterized models, noisy labeled data tends to be classified alongside its genuine counterparts. Our approach suggests that the double descent phenomenon is a consequence of deep learning models learning how to separate the noise from the information in the training set.
@@ 3 Methodology
We first assess the contribution of data noise to the double descent phenomenon. Bartlett et al. (2020) describes an intuitive way of understanding properties of significant over-parameterization in linear regression: fitting the training data exactly but with near-optimal prediction accuracy occurs if and only if there are many low-variance (and hence, unimportant) directions in parameter space where the label noise can be hidden. While non-linear models are different from linear models in terms of parameter space, we propose that neural networks approach optimal prediction through a similar mechanism of concealing label noise. We hypothesize that the emergence of the double descent phenomenon can be attributed to the progressively over-parameterized models effectively isolating noisy data within the training set, thus diminishing the influence of interpolating these noisy data points.
Since this proposition is inspired by Nakkiran et al. (2021)’s empirical conclusion, we initiate our experimentation by replicating the double descent phenomenon as per the experimental framework outlined by this paper. This approach offers a comprehensively manageable setup and establishes a comparative analysis of the phenomenon across diverse tasks and optimization techniques. With an increasing ratio of label noise introduced to the training dataset, the test error peak at the interpolation threshold becomes evident, signifying a definite correlation with the pronounced double descent phenomenon.
@@ 3.1 Experiment Setup
  • Standard Convolutional Networks (CNNs): We consider a family of standard convolutional neural networks formed by 4 convolutional stages of controlled base width [$k$, $2k$, $4k$, $8k$], for $k$ in the range of [1, 64], along with a fully connected layer as the classifier. The MaxPool is [2, 2, 2, 4]. For all the convolution layers, the kernel size = 3, stride = 1, and padding = 1. This architecture implementation is adopted from Nakkiran et al. (2021).
@@
All networks are trained using the Stochastic Gradient Descent (SGD) optimizer with Cross-Entropy Loss, zero momentum, and decreasing learning rate. A batch size of 128 is employed, and no explicit data augmentation or dropout techniques are utilized in our experimental setup. Utilizing a sequence of models trained under a uniform training scheme showcases the double descent phenomenon across different models. We refer our reader to Appendix A for further information regarding our experiment setup and scaling of model parameterization with layer width k.
@@ 3.2 Introduction of Label Noise
While the presence of label errors has been established in the majority of commonly used benchmarking datasets including MNIST and CIFAR-10 (Zhang, 2017; Northcutt et al., 2021b), for the purpose of facilitating comparative analysis of our findings, we intentionally introduce explicit label noise. In our experimental setup, label noise with probability $p$ signifies that during training, samples are assigned a uniformly random label with probability $p$, while possessing the correct label with probability $(1-p)$ otherwise. We interpret a neural network with input variable $x\in R^{d}$ and learnable parameter $\theta$, incorporating all weights and biases. The training dataset is characterized as a collection of training points $(x_{n},y_{n})$, where $n$ ranges from 1 to $N$. In a training dataset with $N$ data points, it comprises $m=p\times N$ noise labeled data points and $\hat{m}=(1-p)\times N$ clean data points.
@@
It is important to note that the label noise is sampled only once and not per epoch, and this original and random label information is stored before each experiment. Each experiment is replicated multiple times, and the final result is obtained by averaging the outcomes. Owing to the presence of inherent label errors, the probability $p$ is solely employed to signify the extent of explicit label noise and does not reflect the overall noise level present in the dataset.
3.3 Learned Feature Space Inference Strategy
Our primary research inquiry aims to gain insights into the diverse strategies employed by deep neural networks of various widths in interpolating noisy labeled data in classification tasks. Based on the existence of benign over-parameterization, we assume that test images, akin to training images mislabeled, are more likely to be correctly classified by over-parameterized models. For instance, we anticipate that an optimal classifier should yield accurate predictions for unseen images, even if it was trained on similar images with an adversarial label. Thus, we further hypothesize that the closest neighbors of mislabeled training images are correctly classified. To validate our hypothesis, we adopted the $k$-nearest neighbour algorithm with cosine similarity to infer the relative locations of clean and noisy labeled data in the learned feature space. This methodology allows us to delineate the prediction strategy through an indirect approach and subsequently compare our findings with the overall generalization performance of these pre-trained models.
@@ 3.3 Learned Feature Space Inference Strategy
We consider L-layer neural networks, while representations on each layer are represented by $f_{l}(X)$, for $l=[1,...,L]$. Our investigation centered on the penultimate layer representations before classification denoted as $f_{L-1}(X)$. Acknowledging the difficulties of inferring high-dimensional feature spaces, our approach mainly entails characterizing the relative positioning of training data points. This involves establishing a connection between the hidden features of noisy data points in the learned feature space and their $k$-nearest neighbors within the complementary subset of clean data. So we calculate the prediction accuracy $P$ of mislabeled training data and the majority of its nearest neighbours in feature space are in the same class:
$P=\frac{\sum^{m}_{i=1}[y_{i}=M(f_{L-1}(x_{i,1}),\ldots,f_{L-1}(x_{i,k}))]}{m},$ (1)
In this context, $(x_{1},y_{1}),\ldots,(x_{m},y_{m})$ denotes the subset of noisy training data with original labels, and $(x_{i,1},y_{i,1}),\ldots,(x_{i,\hat{m}},y_{i,\hat{m}}))$ denotes the subset of clean training data as neighbours to $x_{i}$ for all $i=[1,\ldots,m]$. Given the cosine similarity $S_{C}(A,B):=\frac{A\cdot B}{\|A\|\|B\|}$ between two representations A and B, let $f_{L-1}(x_{i,1}),\ldots,f_{L-1}(x_{i,\hat{m}})$ be a reordering of the representations of the clean training data such that $S_{C}(f_{L-1}(x_{i}),f_{L-1}(x_{i,1}))\leq\cdots\leq S_{C}(f_{L-1}(x_{i}),f_{L-1}(x_{i,\hat{m}}))$. The closest $k$ neighbours are selected, and a Majority function $M(f_{L-1}(x_{i,1}),\ldots,f_{L-1}(x_{i,k}))$ describes the procedure for making predictions based on the majority label among the k-nearest neighbors. The predictions are subsequently compared with the original labels of noisy data, and $P$ denotes the proportion of accurate predictions among all the noisy data.
In line with our hypothesis, the prediction accuracy $P$ is indicative of the neural architecture’s interpolation approach toward these noisy labeled data points. Due to the uncertain distribution of data and their representations, the $k$-NN method functions solely as an estimation technique, and our comparison focuses on variations in $P$.
@@
In the next section, we present our empirical study and results of the learned feature space inference for trained neural networks of various widths in relationship to double descent.